# How to Divide: A Set Partitioning Strategy Balancing the Trade-off Between Intra-Subset Correlation and Inter-Subset Gain Mutual Influence in Distributed Attack Detection Scheduling Task

### v1

@@ IV Simulation Results
$Q_{3,o}=\left[\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 1&1&1&1\\ 0&0&0&0\end{matrix}\right],\quad Q_{4,o}=\left[\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 1&1&1&1\end{matrix}\right].$
@@
According to Kalman decomposition [29], by first deinfluence and then reconstruction, the estimated error $\Delta\hat{x}_{j}^{re}$ of each sensor $i$ and its neighbor sensor $j\in\mathcal{N}_{i}$, which only contains the observable part, can be obtained. The augmented error matrix $\varLambda_{i,k}$ in equation (6) is obtained by summarizing the $\Delta\hat{x}_{j}^{re}$ of all neighbors $j$. It should be noted that the number of non-zero rows of $\Delta\hat{x}_{j}^{re}$ and $Q_{j,o}$ is the same.
Next, we compare the distribution of inter-subset gain mutual influence and intra-subset correlation under different partitioning strategies. Three partitioning strategies are considered: the partitioning strategy based on minimizing inter-subset gain mutual influence (Partitioning Strategy $1$), the suboptimal sensor partitioning strategy based on the Grassmann distance (Partitioning Strategy $2$), and the random division of $100$ sensors into $4$ groups (Partitioning Strategy $3$). For Partitioning Strategy $1$, the observable spaces of $C_{1}$ ($C_{2}$) and $C_{3}$ ($C_{4}$) overlap. Consequently, the partitioning result is that subset $1$ includes all sensors with measurement matrices $C_{1}$ and $C_{3}$, while subset $2$ includes all sensors with measurement matrices $C_{2}$ and $C_{4}$. As shown in TABLE I, the inter-subset benefit influence is 00, while the intra-subset correlation is $0.5$, indicating that this strategy minimizes the inter-subset gain mutual influence. For Partitioning Strategy $2$, sensors are fully correlated only when their measurement matrices are identical. Thus, the partitioning result consists of 4 subsets, each containing one type of sensor. As shown in TABLE I, the inter-subset gain mutual influence is $625$, while the intra-subset correlation is $1$, indicating that this strategy maximizes intra-subset correlation. For Partitioning Strategy $3$, by averaging the results of $100$ completely independent partitioning instances, we obtain the results shown in TABLE I, where neither the inter-subset gain mutual influence nor the intra-subset correlation is optimal. Therefore, the proposed suboptimal partitioning strategy achieves the goal of minimizing inter-subset gain mutual influence while maximizing intra-subset correlation.
TABLE I: Inter-subset Interaction and intra-subset Correlation under different Partitioning Strategy
@@
According to Theorem 1, the gain interactions only exist between $C_{2}$ and $C_{4}$, that is, under these two gain update strategies, the error in the distribution ratio vector only appears in subset $2$. The error curves are shown in Fig. 1. To make the error curves smoother and more intuitive, the evaluation metric used is windowed RMSEs (W-RMSEs), with a window size of $10$ moments. It can be observed that the distribution ratio vector error in subset $2$ is bounded, with an average value of only $042.2871e-$ 04 over the entire time period, while the errors in subsets $1$, $3$, and $4$ remain zero throughout, which is consistent with the conclusion of Theorem 1.
@@
Scenario 2: Next, this simulation verifies that the proposed algorithm can ensure the secure state estimation of the system. Consider a complex distributed network scenario in which $500$ sensors (labeled $1-500$) are randomly deployed in a $200$m ×$\times 200$m area. Each sensor has the capability to communicate with sensors within a $30$m radius, forming an undirected graph. Similar to Scenario $1$, each sensor can independently measure the state of the vehicles, and the measurement matrix is classified into the $4$ types given in equation (IV). First, measurement matrices are randomly assigned to the $500$ sensors, and the number of each type of sensor in the neighbor set of each sensor is recorded. The statistical results indicate that, for almost all sensors, their neighbor sets fail to satisfy the requirement stated in Remark 3, which mandates an equal number of sensors observing different state spaces. In this case, based on the set partitioning result, the communication direction between sensors is dynamically adjusted to control the distribution of neighbor sensors of each sensor. Ultimately, the entire network is transformed into a directed graph, ensuring that the neighbor set of each sensor satisfies Remark 3. Therefore, based on the proposed Algorithm 3.1 or Algorithm 3.2, the neighbor set of each sensor can be divided into $4$ subsets.
@@
At each moment $k=1:100$, the attacker randomly selects no more than half of the target sensors from the set of neighboring sensors of each sensor and launches stealthy attacks on their communication links. Therefore, the estimation results transmitted between sensors may be affected by FDIAs. Based on the D-ADS Algorithm, each sensor can preliminarily exclude a suspicious sensor set before fusing the information from neighboring sensors. This simulation compares the security of the system under different cases, including the case without attack detection, the case using the proposed D-ADS algorithm, and the case without any attack. For the no-attack scenario, the mean of the estimation errors of all sensors at each moment is used to calculate the RMSE. In contrast, for the case where attacks exist, considering that the security of the system can be intuitively described by the maximum estimation error, this simulation calculates the RMSE using the maximum estimation error at each moment. Fig. 2 presents the RMSE curves under these cases, where there are three curves in total: the no-attack case (red), the proposed D-ADS algorithm detection (blue), and the case without attack detection (green). It can be seen that the blue curve is significantly lower than the green curve, indicating that the proposed algorithm ensures system security. However, there is still a certain gap between the blue and red curves, as stealthy attacks are challenging to detect accurately. This observation is consistent with the simulation results in Scenario $1$.
V Conclusions
This paper investigates the performance of distributed attack detection scheduling algorithms in large-scale sensor networks. By analyzing the gain mutual influence between sensor subsets, an set partitioning strategy based on Grassmann distance is proposed, effectively reducing inter-subset mutual influence while enhancing intra-subset correlation. Theoretical analysis and simulation results demonstrate that the proposed strategy not only significantly narrows the performance gap between the D-ADS and ADS algorithm but also reduces computational cost. Future research can explore more advanced feature extraction methods to better capture complex information within the network, thereby further improving the efficiency and accuracy of attack detection.
Appendix A The balance of the attack strategy
@@
Next, we consider the balance of attack intensity. In the literature [29], the average malicious disturbance power of the attack signal $z_{ij}(k)$ is defined as the average of the sum of the squared values of the attack signal $z_{ij}(k)$ on the communication link between sensor $j$ and sensor $i$ over the entire time period. However, since the attacked communication links vary dynamically at each moment, and the same link may experience different attacks at different moment, the definition in [29] is not applicable in this paper.
@@
Based on the previous analysis, if the attacked sensors are completely random at each moment, the expected number of attacked sensors in each subset is $\left\lfloor q_{i}/{m_{i}}\right\rfloor$. Based on Assumption 1, the ratio of the average malicious disturbance power of each subset to the total malicious disturbance power of all attacks is $\left\lfloor q_{i}/m_{i}\right\rfloor/q_{i}=1/m_{i}$. This indicates that the attack intensity in each subset is balanced. This completes the proof. ∎
Then, a basic numerical simulation is utilized to verify that the attacker satisfies Lemma 2. Based on the set partitioning strategy in Algorithm 3.2, the 100 sensors are divided into four subsets: sensors 1–25 form subset 1, sensors 26–50 form subset 2, sensors 51–75 form subset 3, and sensors 76–100 form subset 4. Without considering specific attack types, at each time step, the attacker randomly selects 40 sensors for attack. The number of attacked sensors in each subset is shown in Fig. 3. During the entire time period, the average number of sensors attacked in each subset was 9.83, 9.97, 10.04, and 10.16, respectively, which is close to the ideal average of 10.
Building on this, when specific attack types are considered, the attack intensity distribution for each subset is shown in Fig. 4. The vertical axis represents the ratio of the attack intensity in each subset to the total attack signal intensity, which is ideally 0.25. It can be seen from Fig. 4 that, the intensity ratios fluctuate around 0.25. Specifically, in Fig. 4(a), the mean attack intensity ratios for each subset are 0.2497, 0.2500, 0.2504, and 0.2499, while in Fig. 4(b), the mean ratios are 0.2507, 0.2490, 0.2499, and 0.2504.
@@
Appendix B Distributed Attack Detection Scheduling Algorithm
The ADS algorithm in literature [24] can be extended to the D-ADS algorithm, as shown in Algorithm 3.3. The main difference between the D-ADS and ADS algorithms lies in the sensor selection method, specifically whether it uses centralized selection or distributed two-stage sensor selection. For the $l$-th selection at time step $k$, a suspicious sensor is first chosen from each subset (Stage 1 of sensor selection) and added to the set $\mathcal{S}^{(l)}_{k}$. Then, a sensor is randomly selected from $\mathcal{S}^{(l)}_{k}$ as the result of the $l$-th selection (Stage 2 of sensor selection). The algorithm terminates when $|\mathcal{A}_{i,k}^{(l)}|=q_{i}$.
It should be noted that in steps 4-7, the gains of all remaining sensors $j\in\mathcal{N}_{i}\backslash\mathcal{A}_{i,k}^{(l-1)}$ are updated. However, with the improved partitioning strategy, only the gains of one subset of sensors need to be updated here. The meanings of the parameters in Algorithm 3.3 are summarized in TABLE II [^2] Other variable symbols that are not explained here are the same as literature [24].
@@ Appendix B Distributed Attack Detection Scheduling Algorithm
TABLE II: Summary of Variables in Algorithm 3.3
@@
Next, the performance relationship between the D-ADS algorithm in this paper and the ADS algorithm in literature [24] is analyzed. Both algorithms aim to select the set of suspicious sensors that have the greatest impact on the objective function (6). Therefore, ideally, their performance is consistent, meaning that $\mathcal{A}_{i,k}=\cup^{m_{i}}_{g_{i}=1}{\mathcal{A}_{i,k,g_{i}}}$.
Lemma 3.
For $g_{i}\in\{1,...,m_{i}\}$, $l\in\{0,1,...,q_{i}\}$, define $\delta_{l,g_{i}}$ as $\delta_{l,g_{i}}=\sum_{k=1}^{T}{(}\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}))$ , where $\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}$ represents the set of suspicious sensors after the $l_{g_{i}}$-th selection of the $g_{i}$-th subset at moment $k$. Then, the relationship in literature [24] can be transformed as $\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,g_{i}}-(1-\frac{1}{q_{i}})^{l+1}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}\leq\frac{m_{i}}{q_{i}}\sum_{j=1}^{l+1}(1-\frac{1}{q_{i}})^{l+1-j}B_{i}^{(j)}$. The parameter $B_{i}^{(l)}$ is described in detail in the proof. At this time, the Lemma 3.2 in literature[24] is extended to the distributed case.
Proof.
@@
where the first inequality follows from $f_{k}(\mathcal{A}_{i,k}^{*})\leq\sum^{m_{i}}_{g_{i}=1}f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})})$, the second inequality follows from the submodularity of $f_{k}$, and the third equality follows from the definition of $G_{kj}^{(l+1)}$.
Based on Lemma 3 in literature [30], as well as the fact that $\delta_{l}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ and $\delta_{l,g_{i}}\geq\delta_{l+1,g_{i}}$, we have
@@
Thus, Lemma 1 in literature [24] has been extended to the distributed case. This completes the proof. ∎
Theorem 2.
@@
Essentially, the expectation of the error represents the expected difference between the objective function value of the selected suspicious sensor set and $1-1/e$ times the optimal suspicious sensor set’s objective function value. Therefore, the larger the objective function value of a subset, the greater the upper bound of the error expectation. Consequently, the subset with the largest objective function value has the highest upper bound for error expectation, as shown in equation (B)
@@
where the first inequality follows from literature [25], which states that $\max_{g_{i}}f_{k}(\mathcal{A}_{i,k,g_{i}})\geq\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,k})$, and the second inequality follows from $(1-1/k)^{k}\leq 1/e$.
@@
Based on the Lemma 3 from literature [30] and the definition $\delta_{l}=\sum_{k=1}^{T}\left(f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k}^{(l)})\right)$ for $l=1:q_{i}$, we obtain that, $\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}^{*}\right)}-\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}\right)}+(1-(1-\frac{1}{q_{i}})^{q_{i}})\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,k}^{(0)})-f_{k}(\mathcal{A}_{i,k}^{*}))}\leq\delta_{q_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\delta_{0}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}$, where the inequality $\delta_{q_{i}}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ follows from the diminishing marginal returns property of the submodular function, and the initial condition satisfies $\delta_{0}=\sum_{g_{i}=1}^{m_{i}}\delta_{0,g_{i}}$. Combining equation (28) with the above relation, equation (B) can be rewritten as
@@
Thus, to prove that equation (B) is bounded, it suffices to show that $\mathbb{E}[B_{i}^{(l)}/m_{i}]$ is bounded
@@
where the first inequality follows from the arithmetic mean-geometric mean inequality, and the second follows from Theorem 2 in literature [24].
@@
where $\mathbb{E}[\cdot]$ represents expectation, and $\tilde{\mathcal{O}}$ hides logarithmic terms. This completes the proof. ∎

### v2

@@ 4 Simulation Results
$Q_{3,o}=\left[\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 1&1&1&1\\ 0&0&0&0\end{matrix}\right],\quad Q_{4,o}=\left[\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 1&1&1&1\end{matrix}\right].$ (27)
@@
According to Kalman decomposition (Suo et al., 2024b), by first deinfluence and then reconstruction, the estimated error $\Delta\hat{x}_{j}^{re}$ of each sensor $i$ and its neighbor sensor $j\in\mathcal{N}_{i}$, which only contains the observable part, can be obtained. The augmented error matrix $\varLambda_{i,k}$ in equation (6) is obtained by summarizing the $\Delta\hat{x}_{j}^{re}$ of all neighbors $j$. It should be noted that the number of non-zero rows of $\Delta\hat{x}_{j}^{re}$ and $Q_{j,o}$ is the same.
Next, we compare the distribution of inter-subset gain mutual influence and intra-subset correlation under different partitioning strategies. Three partitioning strategies are considered: the partitioning strategy based on minimizing inter-subset gain mutual influence (Partitioning Strategy $1$), the suboptimal sensor partitioning strategy based on the Grassmann distance (Partitioning Strategy $2$), and the random division of $100$ sensors into $4$ groups (Partitioning Strategy $3$). For Partitioning Strategy $1$, the observable spaces of $C_{1}$ ($C_{2}$) and $C_{3}$ ($C_{4}$) overlap. Consequently, the partitioning result is that subset $1$ includes all sensors with measurement matrices $C_{1}$ and $C_{3}$, while subset $2$ includes all sensors with measurement matrices $C_{2}$ and $C_{4}$. As shown in TABLE 1, the inter-subset benefit influence is 00, while the intra-subset correlation is $0.5$, indicating that this strategy minimizes the inter-subset gain mutual influence. For Partitioning Strategy $2$, sensors are fully correlated only when their measurement matrices are identical. Thus, the partitioning result consists of 4 subsets, each containing one type of sensor. As shown in TABLE 1, the inter-subset gain mutual influence is $625$, while the intra-subset correlation is $1$, indicating that this strategy maximizes intra-subset correlation. For Partitioning Strategy $3$, by averaging the results of $100$ completely independent partitioning instances, we obtain the results shown in TABLE 1, where neither the inter-subset gain mutual influence nor the intra-subset correlation is optimal. Therefore, the proposed suboptimal partitioning strategy achieves the goal of minimizing inter-subset gain mutual influence while maximizing intra-subset correlation.
Table 1: Inter-subset Interaction and intra-subset Correlation under different Partitioning Strategy
@@
According to Theorem 1, the gain interactions only exist between $C_{2}$ and $C_{4}$, that is, under these two gain update strategies, the error in the distribution ratio vector only appears in subset $2$. The error curves are shown in Figure 1. To make the error curves smoother and more intuitive, the evaluation metric used is windowed RMSEs (W-RMSEs), with a window size of $10$ moments. It can be observed that the distribution ratio vector error in subset $2$ is bounded, with an average value of only $042.2871e-$ 04 over the entire time period, while the errors in subsets $1$, $3$, and $4$ remain zero throughout, which is consistent with the conclusion of Theorem 1.
@@
Scenario 2: Next, this simulation verifies that the proposed algorithm can ensure the secure state estimation of the system. Consider a complex distributed network scenario in which $500$ sensors (labeled $1-500$) are randomly deployed in a $200$m ×$\times 200$m area. Each sensor has the capability to communicate with sensors within a $30$m radius, forming an undirected graph. Similar to Scenario $1$, each sensor can independently measure the state of the vehicles, and the measurement matrix is classified into the $4$ types given in equation (4). First, measurement matrices are randomly assigned to the $500$ sensors, and the number of each type of sensor in the neighbor set of each sensor is recorded. The statistical results indicate that, for almost all sensors, their neighbor sets fail to satisfy the requirement stated in Remark 3, which mandates an equal number of sensors observing different state spaces. In this case, based on the set partitioning result, the communication direction between sensors is dynamically adjusted to control the distribution of neighbor sensors of each sensor. Ultimately, the entire network is transformed into a directed graph, ensuring that the neighbor set of each sensor satisfies Remark 3. Therefore, based on the proposed Algorithm 3.1 or Algorithm 3.2, the neighbor set of each sensor can be divided into $4$ subsets.
@@
At each moment $k=1:100$, the attacker randomly selects no more than half of the target sensors from the set of neighboring sensors of each sensor and launches stealthy attacks on their communication links. Therefore, the estimation results transmitted between sensors may be affected by FDIAs. Based on the D-ADS Algorithm, each sensor can preliminarily exclude a suspicious sensor set before fusing the information from neighboring sensors. This simulation compares the security of the system under different cases, including the case without attack detection, the case using the proposed D-ADS algorithm, and the case without any attack. For the no-attack scenario, the mean of the estimation errors of all sensors at each moment is used to calculate the RMSE. In contrast, for the case where attacks exist, considering that the security of the system can be intuitively described by the maximum estimation error, this simulation calculates the RMSE using the maximum estimation error at each moment. Figure 2 presents the RMSE curves under these cases, where there are three curves in total: the no-attack case (red), the proposed D-ADS algorithm detection (blue), and the case without attack detection (green). It can be seen that the blue curve is significantly lower than the green curve, indicating that the proposed algorithm ensures system security. However, there is still a certain gap between the blue and red curves, as stealthy attacks are challenging to detect accurately. This observation is consistent with the simulation results in Scenario $1$.
5 Conclusions
This paper investigates the performance of distributed attack detection scheduling algorithms in large-scale IoT networks. By analyzing the gain mutual influence between sensor subsets, an set partitioning strategy based on Grassmann distance is proposed, effectively reducing inter-subset mutual influence while enhancing intra-subset correlation. Theoretical analysis and simulation results demonstrate that the proposed strategy not only significantly narrows the performance gap between the D-ADS and ADS algorithm but also reduces computational cost. Future research can explore more advanced feature extraction methods to better capture complex information within the network, thereby further improving the efficiency and accuracy of attack detection.
6 APPENDIX A: The balance of the attack strategy
@@
Next, we consider the balance of attack intensity. In the literature Suo et al. (2024b), the average malicious disturbance power of the attack signal $z_{ij}(k)$ is defined as the average of the sum of the squared values of the attack signal $z_{ij}(k)$ on the communication link between sensor $j$ and sensor $i$ over the entire time period. However, since the attacked communication links vary dynamically at each moment, and the same link may experience different attacks at different moment, the definition in Suo et al. (2024b) is not applicable in this paper.
@@
Based on the previous analysis, if the attacked sensors are completely random at each moment, the expected number of attacked sensors in each subset is $\left\lfloor q_{i}/{m_{i}}\right\rfloor$. Based on Assumption 1, the ratio of the average malicious disturbance power of each subset to the total malicious disturbance power of all attacks is $\left\lfloor q_{i}/m_{i}\right\rfloor/q_{i}=1/m_{i}$. This indicates that the attack intensity in each subset is balanced. This completes the proof.
Then, a basic numerical simulation is utilized to verify that the attacker satisfies Lemma 2. Based on the set partitioning strategy in Algorithm 3.2, the 100 sensors are divided into four subsets: sensors 1–25 form subset 1, sensors 26–50 form subset 2, sensors 51–75 form subset 3, and sensors 76–100 form subset 4. Without considering specific attack types, at each time step, the attacker randomly selects 40 sensors for attack. The number of attacked sensors in each subset is shown in Figure 3. During the entire time period, the average number of sensors attacked in each subset was 9.83, 9.97, 10.04, and 10.16, respectively, which is close to the ideal average of 10.
Building on this, when specific attack types are considered, the attack intensity distribution for each subset is shown in Figure 4. The vertical axis represents the ratio of the attack intensity in each subset to the total attack signal intensity, which is ideally 0.25. It can be seen from Figure 4 that, the intensity ratios fluctuate around 0.25. Specifically, in Figure 4(a), the mean attack intensity ratios for each subset are 0.2497, 0.2500, 0.2504, and 0.2499, while in Figure 4(b), the mean ratios are 0.2507, 0.2490, 0.2499, and 0.2504.
@@
7 APPENDIX B: Distributed Attack Detection Scheduling Algorithm
The ADS algorithm in literature Suo et al. (2024c) can be extended to the D-ADS algorithm, as shown in Algorithm 3.3. The main difference between the D-ADS and ADS algorithms lies in the sensor selection method, specifically whether it uses centralized selection or distributed two-stage sensor selection. For the $l$-th selection at time step $k$, a suspicious sensor is first chosen from each subset (Stage 1 of sensor selection) and added to the set $\mathcal{S}^{(l)}_{k}$. Then, a sensor is randomly selected from $\mathcal{S}^{(l)}_{k}$ as the result of the $l$-th selection (Stage 2 of sensor selection). The algorithm terminates when $|\mathcal{A}_{i,k}^{(l)}|=q_{i}$.
It should be noted that in steps 4-7, the gains of all remaining sensors $j\in\mathcal{N}_{i}\backslash\mathcal{A}_{i,k}^{(l-1)}$ are updated. However, with the improved partitioning strategy, only the gains of one subset of sensors need to be updated here. The meanings of the parameters in Algorithm 3.3 are summarized in TABLE $2$Other variable symbols that are not explained here are the same as literature Suo et al. (2024c).
@@ 7 APPENDIX B: Distributed Attack Detection Scheduling Algorithm
Table 2: Summary of Variables in Algorithm 3.3
@@
Next, the performance relationship between the D-ADS algorithm in this paper and the ADS algorithm in literature Suo et al. (2024c) is analyzed. Both algorithms aim to select the set of suspicious sensors that have the greatest impact on the objective function (6). Therefore, ideally, their performance is consistent, meaning that $\mathcal{A}_{i,k}=\cup^{m_{i}}_{g_{i}=1}{\mathcal{A}_{i,k,g_{i}}}$.
Lemma 3
For $g_{i}\in\{1,...,m_{i}\}$, $l\in\{0,1,...,q_{i}\}$, define $\delta_{l,g_{i}}$ as $\delta_{l,g_{i}}=\sum_{k=1}^{T}{(}\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}))$ , where $\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}$ represents the set of suspicious sensors after the $l_{g_{i}}$-th selection of the $g_{i}$-th subset at moment $k$. Then, the relationship in literature Suo et al. (2024c) can be transformed as $\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,g_{i}}-(1-\frac{1}{q_{i}})^{l+1}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}\leq\frac{m_{i}}{q_{i}}\sum_{j=1}^{l+1}(1-\frac{1}{q_{i}})^{l+1-j}B_{i}^{(j)}$. The parameter $B_{i}^{(l)}$ is described in detail in the proof. At this time, the Lemma 3.2 in literatureSuo et al. (2024c) is extended to the distributed case.
{pf}
@@
!where the first inequality follows from $f_{k}(\mathcal{A}_{i,k}^{*})\leq\sum^{m_{i}}_{g_{i}=1}f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})})$, the second inequality follows from the submodularity of $f_{k}$, and the third equality follows from the definition of $G_{kj}^{(l+1)}$.
Based on Lemma 3 in literature Matsuoka et al. (2021), as well as the fact that $\delta_{l}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ and $\delta_{l,g_{i}}\geq\delta_{l+1,g_{i}}$, we have
@@
Thus, Lemma 1 in literature Suo et al. (2024c) has been extended to the distributed case. This completes the proof.
Theorem 2
@@
Essentially, the expectation of the error represents the expected difference between the objective function value of the selected suspicious sensor set and $1-1/e$ times the optimal suspicious sensor set’s objective function value. Therefore, the larger the objective function value of a subset, the greater the upper bound of the error expectation. Consequently, the subset with the largest objective function value has the highest upper bound for error expectation, as shown in equation (7)
@@
where the first inequality follows from literature Mirzasoleiman et al. (2016), which states that $\max_{g_{i}}f_{k}(\mathcal{A}_{i,k,g_{i}})\geq\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,k})$, and the second inequality follows from $(1-1/k)^{k}\leq 1/e$.
@@
Based on the Lemma 3 from literature Matsuoka et al. (2021) and the definition $\delta_{l}=\sum_{k=1}^{T}\left(f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k}^{(l)})\right)$ for $l=1:q_{i}$, we obtain that, $\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}^{*}\right)}-\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}\right)}+(1-(1-\frac{1}{q_{i}})^{q_{i}})\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,k}^{(0)})-f_{k}(\mathcal{A}_{i,k}^{*}))}\leq\delta_{q_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\delta_{0}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}$, where the inequality $\delta_{q_{i}}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ follows from the diminishing marginal returns property of the submodular function, and the initial condition satisfies $\delta_{0}=\sum_{g_{i}=1}^{m_{i}}\delta_{0,g_{i}}$. Combining equation (30) with the above relation, equation (7) can be rewritten as
@@
Thus, to prove that equation (7) is bounded, it suffices to show that $\mathbb{E}[B_{i}^{(l)}/m_{i}]$ is bounded
@@
where the first inequality follows from the arithmetic mean-geometric mean inequality, and the second follows from Theorem 2 in literature Suo et al. (2024c).
@@
where $\mathbb{E}[\cdot]$ represents expectation, and $\tilde{\mathcal{O}}$ hides logarithmic terms. This completes the proof.
[…]
Declaration of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. Generative AI and AI-assisted technologies are only applied to the writing process to improve the readability and language of the manuscript.